Given measure of angle X is 30 degrees, measure of angle Y is 30 degrees, line wz is 10 and line xy equals 18. Find the area of trapezoid wxyz
Answer by Boreal(15235)

This is an isosceles trapezoid
Lengths of bases are 10,18
This is difficult to show without a picture, bu the altitude is between w and XY line and z and xy line, and it intersects 4 units from each end, because the trapezoid is symmetrical.
This is a 30-60-90 triangle, where the side opposite the 30 is the altitude. The adjacent side is 4, so the opposite side is (4/3) sqrt (3). The hypotenuse, which you don't need here, happens to be (8/3) sqrt (3).
A= (1/2) h (b1+b2)
=(1/2) (4/3) sqrt (3)*28
=14 *(4/3 sqrt (3))
=(56/3) sqrt (3)
=32.33 square units.